The number G of gears a machine can make varies directly as the time T it operates. If it can make 3825 gears in 4 hours, how many gears can it make in 7 hours?
Answer by checkley75(3666)

4/3825=7/X NOW CROSS MULTIPLY
4X=7*3825
4X=26775
X=26775/4
X=6693.75 GEARS CAN BE MADE IN 7 HOURS.